Delph Company uses a job-order costing system and has two manufacturing departments—Molding and Fabrication. The company provided the following estimates at the beginning of the year: Molding Fabrication Total Machine-hours 25,000 35,000 60,000 Fixed manufacturing overhead costs $ 730,000 $ 280,000 $ 1,010,000 Variable manufacturing overhead cost per machine-hour $ 5.80 $ 5.80 During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-70 and Job C-200. It provided the following information related to those two jobs: Job D-70: Molding Fabrication Total Direct materials cost $ 373,000 $ 328,000 $ 701,000 Direct labor cost $ 220,000 $ 150,000 $ 370,000 Machine-hours 17,000 8,000 25,000 Job C-200: Molding Fabrication Total Direct materials cost $ 300,000 $ 250,000 $ 550,000 Direct labor cost $ 180,000 $ 280,000 $ 460,000 Machine-hours 8,000 27,000 35,000 Delph had no underapplied or overapplied manufacturing overhead during the year. Required: 1. Assume Delph uses a plantwide predetermined overhead rate

based on machine-hours. a. Compute the plantwide predetermined overhead rate. b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200. c. If Delph establishes bid prices that are 140% of total manufacturing cost, what bid prices would it have established for Job D-70 and Job C-200? d. What is Delph’s cost of goods sold for the year?
Answer #1
1a
Estimated Variable manufacturing overhead 348000 =(25000*5.80)+(35000*5.80)
Estimated Fixed manufacturing overhead 1010000
Total Estimated manufacturing overhead 1358000
Divide by Estimated Machine hours 60000
Predetermined overhead rate 22.63 per MH
1b
Job D-70 Job C-200
Direct materials cost 701000 550000
Direct labor cost 370000 460000
Overhead applied 565750 792050
Total manufacturing cost 1636750 1802050
1c
Job D-70 Job C-200
Total manufacturing cost 1636750 1802050
Add: Markup @40% 654700 720820
Bid prices 2291450 2522870
1d
Total manufacturing cost Job D-70 1636750
Total manufacturing cost Job C-200 1802050
Cost of goods sold 3438800
Workings:
Overhead applied:
Job D-70 = 25000*22.63= 565750
Job C-200 = 35000*22.63= 792050